# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ 5 Simulations
We demonstrate the performance of our CPT-PG algorithm across three settings: (a) traffic control to show the influence of the probability distortion function, (b) scalability to larger state spaces compared to existing zeroth-order methods, and (c) continuous state-action space in an electricity management application. See appendix I and Table 2 for additional details and simulations in Appendix I.2-I.3 (theory illustrations), I.7 (financial trading) and I.8 (control).
Figure 1: Environments: (Left) Traffic control: red roads in the city center are prone to congestion. (Right) Electricity management: arrows refer to electricity flow.
Figure 2: Returns along the iterations of our PG algorithm for (CPT-PO) for: (Left) different utility functions with the same distortion $w$ in the grid environment, (Center) traffic control. Shaded areas indicate a range of $\pm$ one standard deviation over 20 runs. (Right) Density of the empirical returns obtained by deploying different trained PG policies (from different initializations) for electricity management, density computed using 10,000 runs for each curve. See Appendix I for details.
(a) Traffic control over a grid. We simulate a car agent navigating a city grid, where central roads are faster but risk higher delays (see fig. 1 center). The agent must balance speed against risk by avoiding the city center. In risk-neutral settings, the agent favors faster routes, while risk-averse policies avoid the risky central roads. Fig. 1 (center) demonstrates that our algorithm successfully adapts to different risk-weighted objectives. We also consider solving (CPT-PO) with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. Fig. 2 show the corresponding different CPT returns observed. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b) in appendix I.4.
(b) Scalability to larger state spaces. We compare our PG algorithm with the zeroth-order (CPT-SPSA-G) of L.A. et al. (2016) on MDPs with increasing size $n\times n$. Fig. 3 highlights our algorithm’s scalability, while CPT-SPSA-G suffers with larger grids. See appendix I.4 for details.
@@
Figure 3: Compared performance of our algorithm and CPT-SPSA-G for $n=3,5,9$. The shaded area is a range of $\pm$ one standard deviation over 10 independent runs.
(c) Electricity management. In this application involving continuous state and action spaces, an individual home manages its electricity consumption and storage. We consider a 24-hour time frame where the agent must decide how much electricity to buy/sell, based on solar panel production, market prices, and battery levels. We use public data for selling prices recorded on a national electricity network. We experiment with risk-neutral, risk-averse, and risk-seeking objectives with 3 weight functions $w$. Our algorithm performs well across these scenarios, with the risk-averse policy minimizing downside risk, and the risk-seeking policy maximizing potential gains. Fig. (2) (right) shows the distribution of total returns for different objectives. The risk-averse policy avoids selling a lot of electricity and tends to keep it stored until the end of the day. Conversely, the risk-seeking policy aggressively sells energy when the markets are high at the cost of possibly having to buy it again later in the day. See appendix I.6 for more details.
@@ 6 Related Work
Prospect Theory and its sibling, CPT (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Barberis, 2013), were first integrated with RL by L.A. et al. (2016). Since then, only a few studies have explored the CPT-RL framework (Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Ethayarajh et al., 2024). Notably, Borkar and Chandak (2021) proposed a Q-learning algorithm for CPT-based policy optimization, while Ramasubramanian et al. (2021) developed value-based methods estimating CPT values using dynamic programming. Their approach optimizes a sum of CPT-transformed period costs, making it amenable to dynamic programming (see remark 1 therein). In contrast, our CPT formulation is different: we maximize the CPT value of the return of a policy (CPT-PO). This objective lacks an additive structure, hence does not satisfy a Bellman equation, rendering dynamic programming approaches inapplicable. Additionally, prior value-based methods are limited to finite state-action spaces, whereas our PG algorithm is also suitable for continuous state action settings, as shown in our experiments. More recently, Ethayarajh et al. (2024) incorporated CPT (without probability distortion) for fine-tuning large language models with human feedback. Our work complements prior CPT-RL studies (L.A. et al., 2016; Jie et al., 2018) that rely on zeroth-order SPSA methods (Spall, 1992). Instead, we introduce a PG algorithm that leverages first-order information, exploiting the structure of CPT values applied to cumulative rewards (see Section 4 for further comparison). Unlike existing PG approaches in risk-sensitive RL, our method explicitly accounts for probability distortion and S-shaped utility transformations, key aspects of CPT. For a broader discussion on CPT-RL, convex RL, and risk-sensitive RL, see Appendix C. A diagram summarizing these connections is provided in Appendix F.1.
@@ 7 Conclusion
We explored a CPT-based variant of RL to better model human decision-making, providing new insights into optimal policies and their differences from standard RL. We then introduced a novel PG algorithm for CPT-PO, demonstrating its scalability over prior methods and its effectiveness in simple applications. Our work opens several avenues for future research. CPT relies on predefined utility and distortion functions—can these be learned to better align with human preferences? Additionally, modeling collective human behavior through CPT in incentive design problems could be a promising direction. We hope this work inspires further research into capturing the complexity of human decision-making, beyond what expected utility theory alone can account for.
@@ Acknowledgements
We thank the anonymous reviewers for their valuable comments that have helped improve this work.
Appendix A Notation for Policy Classes
@@
  • $\Pi_{NM}:=\{\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$ is the set of non-Markovian policies,444 By ‘non-Markovian’, we mean ‘non necessarily Markovian’ policies including Markovian ones. Elements of $\Pi_{NM}-\Pi_{M,NS}$ can be designated as ‘stricly non-Markovian’ policies. Likewise, by ‘non stationary’, we mean ‘non necessarily stationary’, and by ‘stochastic’ we mean ‘non necessarily deterministic’.
@@
Appendix B CPT-RL vs Risk-Sensitive RL: An Illustration
@@
  • Rewards: Safe steps incur a small penalty $(-10)$. The risky state (2,4) offers a high reward $(10^{6})$ with low probability $(0.01)$, otherwise a large penalty $(-10^{3})$. The remaining risky states give the same penalty. Reaching the goal provides a reward of $50,000$.
@@
Algorithms for policy optimization. We compare our CPT-PG algorithm with an exponential risk-sensitive PG algorithm (using an exponential utility $(\mathcal{U}(x)=\frac{1}{\beta}\exp(-\beta x)),\beta=0.1$, without probability weighting). For our CPT-PG algorithm, we use Kahneman and Tversky’s utility function with parameters $(x_{0}=1,\lambda=3,\alpha=0.6)$ (see section 2) and a probability weight function which is piecewise affine given by the following coefficients $w=[4,0,0.8,0.2,2.7,-1.7,0.1,0.9]$ where $[a_{1},a_{2},a_{3},b_{1},b_{2},b_{3},c_{1},c_{2}]$ stands for a piecewise affine $w$ function with $w(x)=a_{i}x+b_{i}$ for $c_{i-1}<x<c_{i}$. Note that this weight function inflates low probability events. This is one of the key features of CPT that we illustrate. We use a step size $\alpha=0.01$ for both algorithms. For the policy network, we use a simple two-layer feed-forward neural network with a Leaky ReLU activation in the hidden layer and a softmax output layer for action selection.
@@
Figure 5: Heat maps representing the policies trained using our algorithm CPT-PG and an exponential risk-sensitive PG algorithm. Each cell in the $5\times 5$ grid corresponds to a state, risky states are denoted by the letter ‘R’ in the cell, ‘A’ stands for the initial state and ‘G’ for the goal state, all the other states are considered safe (with a zero probability assigned). The color represents the probability of selecting a risky action at each state. The risky state (2,4) is risky in the sense that with low probability $0.01$ it leads to high reward of $10^{6}$ and a penalty of $-10^{3}$ otherwise. The main observation here is that CPT takes more risk in trying to end up in this risky state (low probability high reward) as you can see with the yellow cell above the risky state. Overall the risk profile is different for both methods. This is partly explained by the fact that CPT inflates low probability events thanks to the probability weight function.
@@
Appendix C Extended Related Work Discussion
C.1 Risk-sensitive RL
@@
There is a rich literature around risk sensitive control and RL that we do not hope to give justice to here. We refer the reader to recent comprehensive surveys on the topic [Garcıa and Fernández, 2015, Prashanth et al., 2022] and the references therein. Let us briefly mention that there exist several approaches to risk sensitive RL. These include formulations such as constrained stochastic optimization to control the tolerance to perturbations and stochastic minmax optimization to model robustness with respect to worst case perturbations for instance. Another approach which is more relevant to our paper discussion consists in regularizing or modifying objective functions. Such modifications are based on considering different statistics of the return deviating from the standard expectation such as the variance or the conditional value at risk (e.g. Tamar et al. [2012], Chow and Ghavamzadeh [2014], Chow et al. [2018]) or even considering the entire distribution of the returns like distributional RL [Bellemare et al., 2023]. Another popular objective modification consists in maximizing an exponential criterion (e.g. Borkar [2002], Noorani et al. [2022]) to obtain robust policies w.r.t noise and perturbations of system parameters or variations in the environment. Noorani et al. [2022] designed a model-free REINFORCE algorithm and an actor-critic variant of the algorithm leveraging an (approximate) multiplicative Bellman equation induced by the exponential objective criterion. Moharrami et al. [2024] recently proposed and analyzed similar PG algorithms for the same exponential objective. Vijayan and LA [2023] introduced a PG algorithm for solving risk-sensitive RL for a class of smooth risk measures including some distortion risk measures and a mean-variance risk measure. Their approach is based on simultaneous perturbation stochastic approximation (SPSA) [Bhatnagar et al., 2013] using zeroth-order information to estimate gradients. Our CPT-PO problem covers several of the aforementioned objectives including smooth distortion risk measures and exponential utility as particular cases (see appendix F for more details).
C.2 Convex RL/RL with General Utilities
@@
In the last few years, convex RL (a.k.a. RL with general utilities) [Hazan et al., 2019, Zhang et al., 2020, Zahavy et al., 2021, Geist et al., 2022] has emerged as a framework to unify several problems of interest such as pure exploration, imitation learning or experiment design. More precisely, this line of research is concerned with maximizing a given functional of the state(-action) occupancy measure w.r.t. a policy. To solve this problem, several policy gradient algorithms have been proposed in the literature [Zhang et al., 2021, Bai et al., 2022, Barakat et al., 2023, 2024]. Mutti et al. [2022b, a, 2023a] challenged the initial problem formulation and proposed a finite trial version of the problem which is closer to practical concerns as it consists in maximizing a functional of the empirical state(-action) distribution rather than its true asymptotic counterpart. The particular case of our CPT policy optimization problem without probability distortion (see (EUT-PO) below) coincides with a particular case of the single trial convex RL problem [Mutti et al., 2023b] in which the function of the empirical visitation measure is a linear functional of the reward function (see appendix F.5 for details). However, our general problem is not a particular case of convex RL which does not account for probability distortions. Furthermore, our utility function is in general nonconvex in our setting (see example in Fig 8) and our policy gradient algorithm is model-free. More recently, De Santi et al. [2024] introduced a global RL problem formulation where rewards are globally defined over trajectories instead of locally over states and used submodular optimization tools to solve the resulting non-additive policy optimization problem. While global RL allows to account for trajectory-level global rewards, it does not take into consideration probability distortions. In addition, their investigation is restricted to the setting where the transition model is known whereas our PG algorithm is model-free.
C.3 Cumulative Prospect Theoretic RL

### v3

@@ 5 Simulations
While our main contributions are theoretical and methodological, we provide simulations to illustrate our findings. Our main goals in this section are: (a) to show how our CPT-PG algorithm produces policies illustrating more nuances in capturing human behavior compared to standard expected utility theory and risk sensitive RL; (b) to show that, as expected, our algorithm scales better to larger state spaces than existing zeroth-order methods in a grid MDP with increasing size; (c) to test our algorithm on a finance application to show the flexibility of CPT-RL.
(a) Nuances of CPT-RL. We test our CPT-PG algorithm and compare it to vanilla PG (vPG) and an exponential risk-sensitive PG (ERS-PG) algorithm on two simple 2-bandit action problem instances. In Environment 1 (Gain Bandit): A safe action guarantees a certain gain of 2 and a risky action gives a reward 5 or 0 each with probability 1/2. Clearly the risky action has a higher expected return of 2.5. In Environment 2 (Loss Bandit): We flip the signs of the rewards. The safe action guarantees a certain negative reward of -2 whereas the risky action yields either -5 or 0 with probability 1/2 each. In this case, the risky action has a smaller expected return of -2.5.
Figure 1: Comparison of our CPT-PG algorithm with vanilla PG (vPG) and exponential risk-sensitive (ERS-PG) on a simple 2-action bandit setting. (Upper fig.) Gain lottery setting: vPG trains a policy picking the risky action whereas CPT-PG and ERS-PG choose the safe one. (Lower fig.) Loss lottery: Only CPT picks the risky action. (Left) Recorded distorted returns, (center) evolution of probability of risky action along training steps, (right) Actions prescribed by trained policies. The shaded area is a range of $\pm$ one standard deviation with 5 independent runs with different seeds.
Setting: We train a softmax policy using CPT-PG, vPG and ERS-PG. Vanilla PG optimizes for the standard expected return (identity for $u$ and $w$), CPT-PG using a KT model with S-shaped utility using parameters ($\alpha=0.6$, $\lambda=2.5$) and an S-shaped probability function $w$ under-weighting the probability 1/2. ERS-PG is run using an exponential (concave) risk sensitive utility $u(x)=\eta^{-1}(1-\exp(-\eta x))$ with $\eta=0.5\,.$ We use the exact same parameters for both environments.
Results and interpretation: As shown in Fig. 1 (right), in the Gain Bandit vPG selects the risky arm (higher expected return), whereas both CPT-PG and ERS-PG choose the safe arm—reflecting human risk aversion over gains. In the Loss Bandit, only CPT-PG “flips” to the risky arm, capturing human risk-seeking over losses. Crucially, CPT-PG does this with the same S-shaped utility and probability-weighting parameters in both settings, whereas the concave ERS-PG objective remains risk-averse throughout. This exact “safe in gains, risky in losses” pattern—known as the reflection effect—is the hallmark of Prospect Theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992) from which our simple example is inspired. Neither expected-utility nor a single-parameter risk-sensitive criterion can reproduce both behaviors simultaneously, but CPT-RL can, using a fixed, psychologically-grounded model. While risk-sensitive RL can, with appropriate parameter tuning, replicate some aspects of human decision-making, it often lacks the asymmetric treatment of gains and losses and the probabilistic distortion that characterize human behavior. Overall our main message is that CPT-RL offers more flexibility in modeling. Note also that CPT captures several risk measures as particular cases using discontinuous weights (see App. G.3G.4).
(b) Robustness to state space size. We compare our PG algorithm with the zeroth-order (CPT-SPSA-G) of L.A. et al. (2016) on MDPs with increasing size $n\times n$. Fig. 2 shows that our CPT-PG algorithm scales better to larger grid sizes than CPT-SPSA-G as expected due to its use of (first-order) gradient information. See App. K.5 for details.
@@
Figure 2: Compared performance of our algorithm and CPT-SPSA-G for $n=3,5,9$. The shaded area is a range of $\pm$ one standard deviation over 10 runs.
(c) Application to finance. The goal is to train RL trading agents using our PG algorithm in the CPT-RL setting using a gym trading environment and data from the Bitcoin USD market. See App. K.8 for more details. We test several utility and probability weighting functions including a risk averse exponential of the form $x\mapsto\frac{1}{\beta}(1-\exp(-\beta x))$ with different values of $\beta$ as well as the KT (Kahneman and Tversky) function with different values of the reference point $x_{0}$ to illustrate its influence. In Figure 3, we make three observations. First, the reference point shifts the values of the achieved CPT returns: The smaller the reference point, the larger are the returns (Fig. 3, left). This is because only values larger than the reference point are perceived as positive returns. This illustrates how the subjective perception of the agent of the returns is taken into account by the model. Second, different values of $\beta$ lead to different trajectories overall which can translate to different levels of risk aversion. In particular, the curves do not match the identity utility case in the first episodes and show more or less risk taken towards optimizing the CPT returns. Third, the exponent $\alpha$ in the utility distorts the function and shifts the returns significantly (Fig. 3, right). Lower values of $\alpha$ lead to higher returns in this setting. This parameter $\alpha$ provides a degree of freedom to model the behavior of the agent as per their perception of the returns. Different values of $\alpha$ modify the curvature of the utility function (w.r.t. the reference point $x_{0}=0$ here) which is concave for gains and convex for losses.
Figure 3: Performance of our PG algorithm on a financial trading application. KT refers to Kahneman and Tversky’s utility function, $x_{0}$ is the reference point used in that utility, exp. refers to exponential and $\alpha$ is the parameter used in the definition of KT’s utility. Shaded areas are interquantile (25-75%) margins and curves report the median values over 10 different runs.
Additional simulations. We provide more simulations demonstrating the applicability of our CPT-PG algorithm in different settings, including illustrations of our theoretical results (App. K.3-K.4) and applications to a traffic control example over a grid (App. K.6), an electricity management example (App. K.7) and a control application (App. K.9).
@@ 6 Related Work
Prospect Theory and its sibling, CPT (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Barberis, 2013), were first integrated with RL by L.A. et al. (2016). Since then, only a few studies have explored the CPT-RL framework (Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Ethayarajh et al., 2024). Notably, Borkar and Chandak (2021) proposed a Q-learning algorithm for CPT-based policy optimization, while Ramasubramanian et al. (2021) developed value-based methods estimating CPT values using dynamic programming. Their approach optimizes a sum of CPT-transformed period costs, making it amenable to dynamic programming (see remark 1 therein). In contrast, our CPT formulation is different: we maximize the CPT value of the return of a policy CPT-PO. This objective lacks an additive structure, hence does not satisfy a Bellman equation, rendering dynamic programming approaches inapplicable. Additionally, prior value-based methods are limited to finite state-action spaces, whereas our PG algorithm is also suitable for continuous state action settings, as shown in our experiments. More recently, Ethayarajh et al. (2024) incorporated CPT (without probability distortion) for fine-tuning large language models with human feedback. Our work complements prior CPT-RL studies (L.A. et al., 2016; Jie et al., 2018) that rely on zeroth-order SPSA methods (Spall, 1992). Instead, we introduce a PG algorithm that leverages first-order information, exploiting the structure of CPT values applied to cumulative rewards (see Section 3 for further comparison). Unlike existing PG approaches in risk-sensitive RL, our method explicitly accounts for probability distortion and S-shaped utility transformations, key aspects of CPT. For a broader discussion on CPT-RL, convex RL, and risk-sensitive RL, see App. D. A diagram summarizing these connections is provided in App. G.2.
@@ 7 Conclusion
We introduced a theoretical policy-gradient framework for CPT-PO, paving the way for more efficient algorithms in practice. Looking ahead, an important direction is learning the utility and probability distortion functions directly from data to better align with human preferences; another is modeling collective human behavior for incentive design via CPT. By advancing behaviorally aligned RL, we aim to stimulate further research on more nuanced, human-aware decision making beyond traditional expected-utility approaches, bridging machine learning, behavioral economics, psychology, and real-world needs.
@@ Acknowledgements
We thank anonymous reviewers for their valuable comments that have helped improve this work.
Appendix A Broader Impact Statement
This paper contributes to the development of human-centered sequential decision-making models incorporating cognitive and psychological biases, which are essential for high-stakes, socially beneficial applications. Despite its importance, this perspective remains underexplored in RL, and our work helps fill this gap. We provide a theoretical framework for CPT-RL policy optimization and test it in simple applications for diverse domains such as energy, traffic control and finance, where improved decision-making can directly benefit society. Looking ahead, we envision that CPT-RL could be valuable in other human-centric fields such as legal and ethical decision-making, cybersecurity, and human-robot interaction, each carrying significant societal implications. Further practical research is required in this direction.
While our primary focus is theoretical and methodological, it is important to acknowledge potential risks, such as the challenge of accurately modeling human preferences or unintended biases introduced by learned CPT functions. Ensuring ethical, interpretable, and responsible deployment of CPT-RL in real-world applications is an essential direction for future research. By advancing behaviorally aligned RL, we hope this work contributes to more nuanced and human-aware decision-making beyond traditional utility-based approaches, helping to bridge the gap between machine learning, psychology, and real-world societal needs.
Appendix B Notation for Policy Classes
@@
  • $\Pi_{NM}:=\{\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$ is the set of non-Markovian policies, [^4] By ‘non-Markovian’, we mean ‘non necessarily Markovian’ policies including Markovian ones. Elements of $\Pi_{NM}-\Pi_{M,NS}$ can be designated as ‘stricly non-Markovian’ policies. Likewise, by ‘non stationary’, we mean ‘non necessarily stationary’, and by ‘stochastic’ we mean ‘non necessarily deterministic’.
@@
Appendix C CPT-RL vs Risk-sensitive RL: An Illustration
@@
  • Rewards: Safe steps incur a small penalty (10)(-10). The risky state (2,4) offers a high reward $(10^{6})$ with low probability $(0.01)$, otherwise a large penalty $(-10^{3})$. The remaining risky states give the same penalty. Reaching the goal provides a reward of $50,000$.
@@
Algorithms for policy optimization. We compare our CPT-PG algorithm with an exponential risk-sensitive PG algorithm (using an exponential utility $(\mathcal{U}(x)=\frac{1}{\beta}\exp(-\beta x)),\beta=0.1$, without probability weighting). For our CPT-PG algorithm, we use Kahneman and Tversky’s utility function with parameters $(x_{0}=1,\lambda=3,\alpha=0.6)$ (see section 2) and a probability weight function which is piecewise affine given by the following coefficients w=[4,0,0.8,0.2,2.7,1.7,0.1,0.9]w=[4,0,0.8,0.2,2.7,-1.7,0.1,0.9] where $a_{1},a_{2},a_{3},b_{1},b_{2},b_{3},c_{1},c_{2}]$ stands for a piecewise affine $w$ function with $w(x)=a_{i}x+b_{i}$ for $c_{i-1}<x<c_{i}$. Note that this weight function inflates low probability events. This is one of the key features of CPT that we illustrate. We use a step size $\alpha=0.01$ for both algorithms. For the policy network, we use a simple two-layer feed-forward neural network with a Leaky ReLU activation in the hidden layer and a softmax output layer for action selection.
@@
Figure 5: Heat maps representing the policies trained using our algorithm CPT-PG and an exponential risk-sensitive PG algorithm. Each cell in the $5\times 5$ grid corresponds to a state, risky states are denoted by the letter ‘R’ in the cell, ‘A’ stands for the initial state and ‘G’ for the goal state, all the other states are considered safe (with a zero probability assigned). The color represents the probability of selecting a risky action at each state. The risky state (2,4) is risky in the sense that with low probability $0.01$ it leads to high reward of $10^{6}$ and a penalty of $10^{3}$ otherwise. The main observation here is that CPT takes more risk in trying to end up in this risky state (low probability high reward) as you can see with the yellow cell above the risky state. Overall the risk profile is different for both methods. This is partly explained by the fact that CPT inflates low probability events thanks to the probability weight function.
@@
Appendix D Extended Related Work Discussion
D.1 Risk-sensitive RL
@@
There is a rich literature around risk sensitive control and RL that we do not hope to give justice to here. We refer the reader to recent comprehensive surveys on the topic [Garcıa and Fernández, 2015, Prashanth et al., 2022] and the references therein. Let us briefly mention that there exist several approaches to risk sensitive RL. These include formulations such as constrained stochastic optimization to control the tolerance to perturbations and stochastic minmax optimization to model robustness with respect to worst case perturbations for instance. Another approach which is more relevant to our paper discussion consists in regularizing or modifying objective functions. Such modifications are based on considering different statistics of the return deviating from the standard expectation such as the variance or the conditional value at risk (e.g. Tamar et al. [2012], Chow and Ghavamzadeh [2014], Chow et al. [2018]) or even considering the entire distribution of the returns like distributional RL [Bellemare et al., 2023]. Another popular objective modification consists in maximizing an exponential criterion (e.g. Borkar [2002], Noorani et al. [2022]) to obtain robust policies w.r.t noise and perturbations of system parameters or variations in the environment. Noorani et al. [2022] designed a model-free REINFORCE algorithm and an actor-critic variant of the algorithm leveraging an (approximate) multiplicative Bellman equation induced by the exponential objective criterion. Moharrami et al. [2024] recently proposed and analyzed similar PG algorithms for the same exponential objective. Vijayan and LA [2023] introduced a PG algorithm for solving risk-sensitive RL for a class of smooth risk measures including some distortion risk measures and a mean-variance risk measure. Their approach is based on simultaneous perturbation stochastic approximation (SPSA) [Bhatnagar et al., 2013] using zeroth-order information to estimate gradients. Our CPT-PO problem covers several of the aforementioned objectives including smooth distortion risk measures and exponential utility as particular cases (see App. G for more details).
D.2 Convex RL/RL with General Utilities
@@
In the last few years, convex RL (a.k.a. RL with general utilities) [Hazan et al., 2019, Zhang et al., 2020, Zahavy et al., 2021, Geist et al., 2022] has emerged as a framework to unify several problems of interest such as pure exploration, imitation learning or experiment design. More precisely, this line of research is concerned with maximizing a given functional of the state(-action) occupancy measure w.r.t. a policy. To solve this problem, several policy gradient algorithms have been proposed in the literature [Zhang et al., 2021, Bai et al., 2022, Barakat et al., 2023, 2024]. Mutti et al. [2022b, a, 2023a] challenged the initial problem formulation and proposed a finite trial version of the problem which is closer to practical concerns as it consists in maximizing a functional of the empirical state(-action) distribution rather than its true asymptotic counterpart. The particular case of our CPT policy optimization problem without probability distortion (see EUT-PO below) coincides with a particular case of the single trial convex RL problem [Mutti et al., 2023b] in which the function of the empirical visitation measure is a linear functional of the reward function (see App. G.6 for details). However, our general problem is not a particular case of convex RL which does not account for probability distortions. Furthermore, our utility function is in general nonconvex in our setting (see example in Fig 8) and our policy gradient algorithm is model-free. More recently, De Santi et al. [2024] introduced a global RL problem formulation where rewards are globally defined over trajectories instead of locally over states and used submodular optimization tools to solve the resulting non-additive policy optimization problem. While global RL allows to account for trajectory-level global rewards, it does not take into consideration probability distortions. In addition, their investigation is restricted to the setting where the transition model is known whereas our PG algorithm is model-free.
D.3 Cumulative Prospect Theoretic RL